A scientist mixes water (containing no salt) with a solution that contains
40%
salt. She wants to obtain
200
ounces of a mixture that is
5%
salt. How many ounces of water and how many ounces of the
40%
salt solution should she use

Answer by Alan3354(69443)

A scientist mixes water (containing no salt) with a solution that contains
40% salt. She wants to obtain 200 ounces of a mixture that is 5% salt. How many ounces of water and how many ounces of the 40% salt solution should she use
---------------
W = Water
S = 40% salt (if available)
-------------
W + S = 200
W*0 + S*40 = (W+S)*5
40S = 5W + 5S
5W = 35S
W = 200-S
5(200-S) = 35S
1000 - 5S = 35S
40S = 1000
S = 25